# Online Conformal Probabilistic Numerics via Adaptive Edge-Cloud Offloading

### v2

@@ 1 Introduction
Specifically, the probabilistic linear solver (PLS) (Cockayne et al., 2019a; Hennig et al., 2015; Bartels et al., 2019; Reid et al., 2020; Pförtner et al., 2024) typically treat the solution of a linear system as a Bayesian inference problem, whereby each iteration provides an updated posterior distribution for the solution $x^{*}_{t}$. The posterior distribution reflects the current knowledge of the true solution given the available computing budget. Using this posterior distribution, PLS can construct a highest-probability-density (HPD) set $\mathcal{C}_{t}$ with the aim of covering the true solution $x^{*}_{t}$ with a pre-determined target probability level.
@@
Model misspecification arises from the choice of a distribution that does not reflect the properties of the true solution $x^{*}_{t}$. For example, one typically assumes a Gaussian prior, but the solution $x^{*}_{t}$ may be sparse. This problem is exacerbated if PLS is implemented with prior-dependent search directions, such as BayesCG
(Hegde et al., 2024; Reid et al., 2020; Wenger and Hennig, 2020; Cockayne et al., 2019a, 2022, b), which may reinforce incorrect prior assumptions. Further work on the impact of calibration on PLS includes reference (Reid et al., 2020), which demonstrated that BayesCG search directions yield a slightly optimistic HPD set under the Krylov prior, and references (Hegde et al., 2024; cockayne2021probabilistic), which explored the use of probabilistic stationary iterative methods rather than Bayesian framework.
@@ 3.1 Online Conformal Prediction
Given an arbitrary sequence of input-output pairs $(x_{t},y_{t})\\ \in\mathcal{X}\times\mathcal{Y}$, for $t=1,2,\ldots$, OCP aims to construct a prediction set $\mathcal{C}_{t}(x_{t})$ on the output space $\mathcal{Y}$ that satisfies long-term coverage guarantees. OCP builds on a scoring function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ that measures the extent to which the output $y$ is mismatched to the input $x$. A smaller value $s(x,y)$ indicates that $y$ is a better fit for input $x$. The set $\mathcal{C}_{t}(x_{t})$ is obtained by choosing all the candidate output values $y\in\mathcal{Y}$ with sufficiently small scores $s(x,y)$, i.e.,
@@
If the score function $s(x,y)$ is bounded within the interval $[m,M]$ with $M>m$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$, then the threshold $\lambda_{t}$ can be also proved to be bounded as $\lambda_{t}\in[m-\gamma\alpha,M+\gamma(1-\alpha)]$, where $\lambda_{1}$ is selected within the interval $[m,M]$ (Gibbs and Candes, 2021). Accordingly, the quantity (12) is given as
@@ 3.2 Intermittent Online Conformal Prediction
with $C$ in (12), where the probability is evaluated over the randomness of the Bernoulli random variables
$\text{obs}_{1},\ldots,\text{obs}_{T}$.
@@ 4.1 Score Function
Since matrix $\Sigma_{t}$ in (4) can be expressed as a sequence of rank-$1$ updates to the prior diagonal matrix $\Sigma_{0}$, its eigendecomposition can be efficiently updated using
Given’s rotations (golub2013matrix, Section 12.5). This results in a per-iteration complexity of $\mathcal{O}(n_{t}^{2})$, and an overall complexity of $\mathcal{O}(I_{t}n_{t}^{2})$ after $I_{t}$ updates. Note that the score (20) depends on the specific sequence of search directions (5), although we do not make this dependence explicit in the notation $s(x_{t};A_{t},b_{t})$.
@@ 5.4 Results
Fig. 3(a) demonstrates that both OCP-PLS and OCP-PLS with full cloud-to-edge feedback guarantee the long-term coverage condition (2), thereby validating Proposition 1. In contrast, confirming prior art (Hegde et al., 2024; Wenger and Hennig, 2020), PLS produces an overly conservative HPD set that always covers the true solution. PLS achieves coverage of $1$ by yielding much larger HPD sets than both OCP-PLS and OCP-PLS with full cloud-to-edge feedback, as shown in
Fig. 3(b). Furthermore, as illustrated in (2), OCP-PLS can reduce the cloud-to-edge feedback rate by approximately $40\%$ as compared to OCP-PLS with full cloud-to-edge feedback, while producing sets with similar sizes.
Consider now the setting with a time-varying computing budget at the edge. Fig. 4 marks the times at which the computing budget changes with vertical dashed lines. Fig. 4(a) verifies that both OCP-PLS and OCP-PLS with full cloud-to-edge feedback maintain the guaranteed long-term coverage condition proved in Proposition 1 even under time-varying computing budgets. As in the previous example, PLS provides perfect coverage at the cost of excessively large HPD sets. In contrast, as shown in Fig. 4(b), OCP-PLS responds naturally to changes in edge computing budget, as the set sizes increase during periods of reduced computing budget ($t=1500$ to $t=3500$), while decreasing when the computing budget is higher ($t=3500$ to $t=5000$). Throughout these variations, OCP-PLS consistently
maintains a significantly lower cloud-to-edge feedback rate than OCP-PLS with full cloud-to-edge feedback, as shown in Fig. 4(c), adapting the feedback rate to the varying computing budgets. This demonstrates OCP-PLS’s ability to balance reliability and communication efficiency under dynamic edge computing budgets.
@@ 6 Conclusion
An important challenge in edge-cloud computing architectures is to provide users with timely and reliable responses under limited computational budgets. This work has made steps towards addressing this challenge by leveraging probabilistic numerics (PN) and online conformal prediction (OCP), an online calibration method previously studied in the context of predictive models. The proposed methodology, referred to as OCP-PLS calibrates the HPD sets produced by probabilistic linear solvers (PLS), ensuring provable coverage guarantees for the true solution. OCP-PLS supports an adaptive mechanism for the use of cloud resources that reduces the communication overhead between edge and cloud, as well as the computational load of the cloud processors. Experiments for tasks requiring the sequential solution of linear systems validate the theoretical reliability, as well as the efficiency, of OCP-PLS. Future work may investigate the generalization of the proposed approach to other numerical problems beyond linear systems (Wenger et al., 2022; Pförtner et al., 2022; Hegde et al., 2024).
Appendix A: Extension to Delayed
Cloud-to-Edge Feedback

### v3

@@ 1 Introduction
Specifically, the probabilistic linear solver (PLS) (Cockayne et al., 2019a; Hennig et al., 2015; Bartels et al., 2019; Reid et al., 2023; Pförtner et al., 2024) typically treat the solution of a linear system as a Bayesian inference problem, whereby each iteration provides an updated posterior distribution for the solution $x^{*}_{t}$. The posterior distribution reflects the current knowledge of the true solution given the available computing budget. Using this posterior distribution, PLS can construct a highest-probability-density (HPD) set $\mathcal{C}_{t}$ with the aim of covering the true solution $x^{*}_{t}$ with a pre-determined target probability level.
@@
Model misspecification arises from the choice of a distribution that does not reflect the properties of the true solution $x^{*}_{t}$. For example, one typically assumes a Gaussian prior, but the solution $x^{*}_{t}$ may be sparse. Misspecification may also arise due to simplifications done in the evaluation of the likelihood in order to obtain Gaussian posteriors, as is the case with BayesCG (Hegde et al., 2025; Reid et al., 2023; Wenger and Hennig, 2020; Cockayne et al., 2019a, 2022, b), which may reinforce incorrect prior assumptions. Further work on the impact of calibration on PLS includes reference (Reid et al., 2023), which demonstrated that BayesCG search directions yield a slightly optimistic HPD set under the Krylov prior, and references (Hegde et al., 2025; Cockayne et al., 2021), which explored the use of probabilistic stationary iterative methods rather than Bayesian framework.
@@ 3.1 Online Conformal Prediction
Given an arbitrary sequence of input-output pairs $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$, for $t=1,2,\ldots$, OCP aims to construct a prediction set $\mathcal{C}_{t}(x_{t})$ on the output space $\mathcal{Y}$ that satisfies long-term coverage guarantees. OCP builds on a scoring function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ that measures the extent to which the output $y$ is mismatched to the input $x$. A smaller value $s(x,y)$ indicates that $y$ is a better fit for input $x$. The set $\mathcal{C}_{t}(x_{t})$ is obtained by choosing all the candidate output values $y\in\mathcal{Y}$ with sufficiently small scores $s(x,y)$, i.e.,
@@
If the score function $s(x,y)$ is bounded within the interval $[m,M]$ with $M>m$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$, then the threshold $\lambda_{t}$ can be also proved to be bounded as $\lambda_{t}\in[m-\gamma\alpha,M+\gamma(1-\alpha)]$, where $\lambda_{1}$ is selected within the interval $[m,M]$ (Gibbs and Candes, 2021; Feldman et al., 2022). Accordingly, the quantity (12) is given as
@@ 3.2 Intermittent Online Conformal Prediction
with $C$ in (12), where the probability is evaluated over the randomness of the Bernoulli random variables $\text{obs}_{1},\ldots,\text{obs}_{T}$.
@@ 4.1 Score Function
Since matrix $\Sigma_{t}$ in (4) can be expressed as a sequence of rank-$1$ updates to the prior diagonal matrix $\Sigma_{0}$, its eigendecomposition can be efficiently updated using Given’s rotations (Golub and Van Loan, 2013, Section 12.5). This results in a per-iteration complexity of $\mathcal{O}(n_{t}^{2})$, and an overall complexity of $\mathcal{O}(I_{t}n_{t}^{2})$ after $I_{t}$ updates. Note that the score (20) depends on the specific sequence of search directions (5), although we do not make this dependence explicit in the notation $s(x_{t};A_{t},b_{t})$.
@@ 5.4 Results
Fig. 3(a) demonstrates that both OCP-PLS and OCP-PLS with full cloud-to-edge feedback guarantee the long-term coverage condition (2), thereby validating Proposition 1. In contrast, confirming prior art (Hegde et al., 2025; Wenger and Hennig, 2020), PLS produces an overly conservative HPD set that always covers the true solution. PLS achieves coverage of $1$ by yielding much larger HPD sets than both OCP-PLS and OCP-PLS with full cloud-to-edge feedback, as shown inFig. 3(b). Furthermore, as illustrated in (2), OCP-PLS can reduce the cloud-to-edge feedback rate by approximately $40\%$ as compared to OCP-PLS with full cloud-to-edge feedback, while producing sets with similar sizes.
@@
Consider now the setting with a time-varying computing budget at the edge. Fig. 4 marks the times at which the computing budget changes with vertical dashed lines. Fig. 4(a) verifies that both OCP-PLS and OCP-PLS with full cloud-to-edge feedback maintain the guaranteed long-term coverage condition proved in Proposition 1 even under time-varying computing budgets. As in the previous example, PLS provides perfect coverage at the cost of excessively large HPD sets. In contrast, as shown in Fig. 4(b), OCP-PLS responds naturally to changes in edge computing budget, as the set sizes increase during periods of reduced computing budget ($t=1500$ to $t=3500$), while decreasing when the computing budget is higher ($t=3500$ to $t=5000$). Throughout these variations, OCP-PLS consistently maintains a significantly lower cloud-to-edge feedback rate than OCP-PLS with full cloud-to-edge feedback, as shown in Fig. 4(c), adapting the feedback rate to the varying computing budgets. This demonstrates OCP-PLS’s ability to balance reliability and communication efficiency under dynamic edge computing budgets.
@@ 6 Conclusion
An important challenge in edge-cloud computing architectures is to provide users with timely and reliable responses under limited computational budgets. This work has made steps towards addressing this challenge by leveraging probabilistic numerics (PN) and online conformal prediction (OCP), an online calibration method previously studied in the context of predictive models. The proposed methodology, referred to as OCP-PLS calibrates the HPD sets produced by probabilistic linear solvers (PLS), ensuring provable coverage guarantees for the true solution. OCP-PLS supports an adaptive mechanism for the use of cloud resources that reduces the communication overhead between edge and cloud, as well as the computational load of the cloud processors. Experiments for tasks requiring the sequential solution of linear systems validate the theoretical reliability, as well as the efficiency, of OCP-PLS. Future work may investigate the generalization of the proposed approach to other numerical problems beyond linear systems (Wenger et al., 2022; Pförtner et al., 2022; Hegde et al., 2025).
Appendix A: Extension to Delayed Cloud-to-Edge Feedback